# Hyperspherical Normalization for Scalable Deep Reinforcement Learning

### v1

@@ Appendix G Training Stability
Results. Figure 8 shows the tracked metrics over 1 million training steps. Certain features (e.g., logits) and parameters (e.g., scalers and interpolation vectors) may occasionally exceed unit norm, the overall parameter norms are tightly controlled (Figure 8.(b)-(c)), and gradient magnitudes are consistently balanced across modules (Figure 8.(d)). This leads to the consistent trend and scales of their ELRs over time. We hypothesize that this stable behavior contributes to underpins SimbaV2’s improved performance and scalability.
@@
Appendix H Scalability Effect of Hyperspherical Normalization
@@ I.5 Humanoid Bench
Task PPO SAC TD3 TD3+OFE TQC CrossQ BRO
Ant-v4 1584 [1360, 1815] 3681 [3175, 4187] 3942 [2912, 4972] 7398 [7280, 7516] 3582 [2489, 4675] 6980 [6834, 7126] 7027 [6710, 7343]
HalfCheetah-v4 1744 [1523, 2118] 10484 [9825, 11143] 10574 [9677, 11471] 13758 [13214, 14302] 12349 [11471, 13227] 12893 [11771, 14015] 13747 [12621, 14873]
Hopper-v4 3022 [2633, 3339] 2785 [2151, 3419] 3226 [2911, 3541] 3121 [2615, 3627] 3526 [3302, 3750] 2467 [1855, 3079] 2122 [1655, 2588]
Humanoid-v4 477 [436, 518] 4090 [4545, 5273] 5165 [5020, 5310] 6032 [5698, 6366] 6029 [5498, 6560] 10480 [10307, 10653] 4757 [3139, 6376]
Walker2d-v4 2487 [1907, 3022] 4314 [4058, 4570] 3946 [3654, 4238] 5195 [4683, 5707] 5321 [4999, 5643] 6257 [5277, 7237] 3432 [2064, 4801]
IQM 0.410 [0.110, 0.834] 0.930 [0.812, 1.059] 1.000 [1.000, 1.000] 1.261 [1.035, 1.680] 1.143 [0.971, 1.290] 1.565 [1.394, 1.710] 1.071 [0.828, 1.333]
Median 0.412 [0.071, 0.936] 0.935 [0.787, 1.093] 1.000 [1.000, 1.000] 1.293 [0.967, 1.861] 1.163 [0.910, 1.349] 1.489 [1.317, 1.643] 1.071 [0.884, 1.322]
Mean 0.447 [0.186, 0.725] 0.934 [0.846, 1.027] 1.000 [1.000, 1.000] 1.322 [1.090, 1.615] 1.137 [1.012, 1.261] 1.475 [1.330, 1.608] 1.101 [0.927, 1.278]
@@
Task PPO SAC SR-SAC iQRL BRO MAD-TD
dog-run 26 [26, 28] 36 [3, 69] 6 [5, 8] 380 [336, 424] 374 [338, 411] 437 [396, 478]
dog-stand 129 [122, 139] 102 [39, 164] 24 [17, 31] 926 [897, 955] 966 [956, 977] 967 [952, 982]
dog-trot 31 [30, 34] 29 [6, 52] 13 [8, 18] 713 [516, 909] 783 [717, 848] 867 [805, 929]
dog-walk 40 [37, 43] 38 [11, 65] 8 [7, 9] 866 [827, 905] 931 [920, 942] 924 [906, 943]
humanoid-run 0 [1, 1] 116 [89, 144] 121 [21, 221] 188 [167, 210] 204 [186, 223] 200 [180, 220]
humanoid-stand 5 [5, 6] 352 [225, 479] 161 [-36, 359] 727 [655, 799] 920 [909, 931] 870 [840, 901]
humanoid-walk 1 [1, 2] 273 [128, 418] 122 [-115, 359] 688 [642, 735] 672 [619, 725] 684 [609, 759]
IQM 0.021 [0.003, 0.069] 0.069 [0.042, 0.114] 0.011 [0.007, 0.036] 0.694 [0.528, 0.805] 0.772 [0.662, 0.854] 0.787 [0.691, 0.865]
Median 0.026 [0.001, 0.040] 0.159 [0.08, 0.183] 0.049 [0.01, 0.095] 0.64 [0.516, 0.766] 0.694 [0.615, 0.774] 0.707 [0.634, 0.786]
Mean 0.033 [0.009, 0.068] 0.136 [0.098, 0.175] 0.065 [0.025, 0.114] 0.642 [0.531, 0.747] 0.693 [0.625, 0.757] 0.708 [0.642, 0.771]
@@
Task DreamerV3 TD7 TD-MPC2 CQN-AS Simba SimbaV2
h1-pole-v0 41 [28, 54] 441 [320, 563] 744 [609, 879] 102 [94, 111] 716 [667, 765] 791 [785, 797]
h1-slide-v0 11 [7, 15] 39 [26, 53] 334 [304, 364] 50 [47, 54] 277 [252, 303] 487 [404, 571]
h1-stair-v0 7 [2, 12] 52 [31, 74] 378 [108, 648] 47 [39, 54] 269 [153, 385] 493 [467, 518]
h1-balance-hard-v0 11 [7, 15] 79 [51, 107] 31 [5, 56] 46 [44, 49] 75 [71, 80] 143 [128, 157]
h1-balance-simple-v0 9 [6, 12] 69 [58, 80] 42 [14, 70] 63 [57, 70] 337 [193, 482] 723 [651, 795]
h1-sit-hard-v0 15 [-4, 35] 235 [154, 315] 723 [660, 786] 220 [143, 297] 512 [354, 670] 679 [548, 811]
h1-sit-simple-v0 19 [9, 28] 874 [869, 879] 790 [772, 809] 692 [512, 872] 833 [814, 853] 875 [870, 880]
h1-maze-v0 113 [107, 118] 147 [137, 156] 244 [106, 383] 126 [111, 141] 354 [342, 366] 313 [287, 340]
h1-crawl-v0 248 [176, 319] 582 [563, 600] 962 [959, 965] 564 [544, 585] 923 [904, 942] 946 [933, 959]
h1-hurdle-v0 4 [3, 5] 60 [18, 102] 387 [254, 519] 23 [12, 33] 175 [150, 201] 202 [167, 236]
h1-reach-v0 3203 [2824, 3581] 1409 [998, 1821] 2654 [1951, 3357] 1850 [1232, 2469] 3874 [3220, 4527] 3850 [3272, 4427]
h1-run-v0 4 [2, 6] 91 [54, 128] 778 [763, 793] 22 [9, 34] 232 [185, 279] 415 [307, 524]
h1-stand-v0 15 [7, 22] 433 [138, 727] 798 [779, 817] 439 [213, 665] 772 [701, 843] 814 [770, 857]
h1-walk-v0 8 [1, 16] 33 [22, 45] 814 [813, 815] 56 [37, 74] 550 [391, 709] 845 [840, 850]
IQM 0.007 [0.004, 0.012] 0.134 [0.088, 0.245] 0.734 [0.510, 0.936] 0.094 [0.070, 0.161] 0.521 [0.413, 0.633] 0.799 [0.686, 0.908]
Median 0.021 [0.000, 0.047] 0.284 [0.183, 0.392] 0.696 [0.536, 0.881] 0.226 [0.139, 0.307] 0.598 [0.514, 0.692] 0.781 [0.693, 0.865]
Mean 0.022 [0.000, 0.046] 0.289 [0.207, 0.375] 0.710 [0.562, 0.858] 0.221 [0.157, 0.296] 0.606 [0.536, 0.678] 0.776 [0.705, 0.849]

### v2

@@ Appendix G Training Stability
Results. Figure 9 shows the tracked metrics over 1 million training steps. Certain features (e.g., logits) and parameters (e.g., scalers and interpolation vectors) may occasionally exceed unit norm, the overall parameter norms are tightly controlled (Figure 9.(b)-(c)), and gradient magnitudes are consistently balanced across modules (Figure 9.(d)). This leads to the consistent trend and scales of their ELRs over time. We hypothesize that this stable behavior contributes to SimbaV2’s improved performance and scalability.
@@
Appendix H Additional Experiments
This section complements Section 5.2 by presenting further experiments probing the properties and robustness of SimbaV2:
  • Scalability Effect of Hyperspherical Normalization (Section H.1). Investigate the necessity of hyperspherical normalization for achieving SimbaV2’s scalability.
  • Effectiveness beyond SAC (Section H.2.) Assess SimbaV2’s broader applicability by substituting SAC with DDPG.
H.1 Scalability Effect of Hyperspherical Normalization
@@
H.2 Effectiveness beyond SAC
In Section 5.2, we observe that replacing the neural network of SAC with SimbaV2 consistently improves the performance of a wide range of domains. To assess the broader applicability and robustness of SimbaV2’s architectural advantages beyond a single algorithm, we conducted additional experiments using Deep Deterministic Policy Gradient (Lillicrap, 2015, DDPG), another widely adopted off-policy algorithm for continuous control.
Experimental Setup. We evaluated SimbaV2 against the original Simba (Lee et al., 2024c) and a standard MLP baseline on two challenging continuous control benchmark suites: DMC-Hard and HBench-Hard. All methods utilized the DDPG algorithm as their underlying learning framework. The MLP architecture we adopted consists of a sequence of linear layers followed by $\mathrm{ReLU}$ non-linearities.
Result. Comparative results are presented in Figure 11. On the DMC-Hard benchmark, SimbaV2 achieved performance competitive with the original Simba, with both significantly outperforming the MLP baseline. More notably, on the more complex HBench-Hard benchmark, SimbaV2 demonstrated a clear improvement over Simba. These results indicate that SimbaV2 not only generalizes to the DDPG algorithm but also exhibits enhanced stability and generalization capabilities in more demanding environments, likely attributable to its refined architecture and regularization mechanisms.
Figure 11: DDPG with SimbaV2. Learning curves of SimbaV2, Simba (Lee et al., 2024c), and the MLP baseline on DMC-Hard and HBench-Hard benchmarks using DDPG (Lillicrap, 2015). SimbaV2 performs competitively with Simba on DMC-Hard, both significantly outperforming the MLP baseline. In the more challenging HBench-Hard, SimbaV2 shows clear improvements over Simba, indicating enhanced stability and generalization beyond SAC.
@@ I.5 Humanoid Bench
Task PPO SAC TD3 TD3+OFE TQC REDQ DroQ CrossQ BRO
Ant-v4 1584 [1360, 1815] 5733 [5316, 6151] 3942 [2912, 4972] 7398 [7280, 7516] 3582 [2489, 4675] 5314 [4539, 6090] 5965 [5560, 6370] 6980 [6834, 7126] 7027 [6710, 7343]
HalfCheetah-v4 1744 [1523, 2118] 11320 [10634, 12007] 10574 [9677, 11471] 13758 [13214, 14302] 12349 [11471, 13227] 11505 [10213, 12798] 11070 [10272, 11867] 12893 [11771, 14015] 13747 [12621, 14873]
Hopper-v4 3022 [2633, 3339] 2787 [2249, 3325] 3226 [2911, 3541] 3121 [2615, 3627] 3526 [3302, 3750] 3299 [2730, 3869] 2797 [2387, 3208] 2467 [1855, 3079] 2122 [1655, 2588]
Humanoid-v4 477 [436, 518] 4825 [3784, 5866] 5165 [5020, 5310] 6032 [5698, 6366] 6029 [5498, 6560] 5278 [5127, 5430] 5380 [5353, 5407] 10480 [10307, 10653] 4757 [3139, 6376]
Walker2d-v4 2487 [1907, 3022] 4536 [4229, 4843] 3946 [3654, 4238] 5195 [4683, 5707] 5321 [4999, 5643] 5228 [4836, 5620] 4781 [4539, 5024] 6257 [5277, 7237] 3432 [2064, 4801]
IQM 0.410 [0.110, 0.834] 1.097 [1.050, 1.155] 1.000 [1.000, 1.000] 1.261 [1.035, 1.680] 1.143 [0.971, 1.290] 1.135 [1.086, 1.194] 1.108 [1.055, 1.170] 1.565 [1.394, 1.710] 1.071 [0.828, 1.333]
Median 0.412 [0.071, 0.936] 1.093 [1.028, 1.180] 1.000 [1.000, 1.000] 1.293 [0.967, 1.861] 1.163 [0.910, 1.349] 1.188 [1.086, 1.241] 1.133 [1.068, 1.199] 1.489 [1.317, 1.643] 1.071 [0.884, 1.322]
Mean 0.447 [0.186, 0.725] 1.092 [1.013, 1.166] 1.000 [1.000, 1.000] 1.322 [1.090, 1.615] 1.137 [1.012, 1.261] 1.160 [1.096, 1.224] 1.134 [1.067, 1.205] 1.475 [1.330, 1.608] 1.101 [0.927, 1.278]
@@
Task PPO SAC iQRL BRO MAD-TD
dog-run 26 [26, 28] 36 [3, 69] 380 [336, 424] 374 [338, 411] 437 [396, 478]
dog-stand 129 [122, 139] 102 [39, 164] 926 [897, 955] 966 [956, 977] 967 [952, 982]
dog-trot 31 [30, 34] 29 [6, 52] 713 [516, 909] 783 [717, 848] 867 [805, 929]
dog-walk 40 [37, 43] 38 [11, 65] 866 [827, 905] 931 [920, 942] 924 [906, 943]
humanoid-run 0 [1, 1] 116 [89, 144] 188 [167, 210] 204 [186, 223] 200 [180, 220]
humanoid-stand 5 [5, 6] 352 [225, 479] 727 [655, 799] 920 [909, 931] 870 [840, 901]
humanoid-walk 1 [1, 2] 273 [128, 418] 688 [642, 735] 672 [619, 725] 684 [609, 759]
IQM 0.021 [0.003, 0.069] 0.069 [0.042, 0.114] 0.694 [0.528, 0.805] 0.772 [0.662, 0.854] 0.787 [0.691, 0.865]
Median 0.026 [0.001, 0.040] 0.159 [0.08, 0.183] 0.64 [0.516, 0.766] 0.694 [0.615, 0.774] 0.707 [0.634, 0.786]
Mean 0.033 [0.009, 0.068] 0.136 [0.098, 0.175] 0.642 [0.531, 0.747] 0.693 [0.625, 0.757] 0.708 [0.642, 0.771]
@@
Task DreamerV3 TD7 TD-MPC2 Simba SimbaV2
h1-pole-v0 41 [28, 54] 441 [320, 563] 744 [609, 879] 716 [667, 765] 791 [785, 797]
h1-slide-v0 11 [7, 15] 39 [26, 53] 334 [304, 364] 277 [252, 303] 487 [404, 571]
h1-stair-v0 7 [2, 12] 52 [31, 74] 378 [108, 648] 269 [153, 385] 493 [467, 518]
h1-balance-hard-v0 11 [7, 15] 79 [51, 107] 31 [5, 56] 75 [71, 80] 143 [128, 157]
h1-balance-simple-v0 9 [6, 12] 69 [58, 80] 42 [14, 70] 337 [193, 482] 723 [651, 795]
h1-sit-hard-v0 15 [-4, 35] 235 [154, 315] 723 [660, 786] 512 [354, 670] 679 [548, 811]
h1-sit-simple-v0 19 [9, 28] 874 [869, 879] 790 [772, 809] 833 [814, 853] 875 [870, 880]
h1-maze-v0 113 [107, 118] 147 [137, 156] 244 [106, 383] 354 [342, 366] 313 [287, 340]
h1-crawl-v0 248 [176, 319] 582 [563, 600] 962 [959, 965] 923 [904, 942] 946 [933, 959]
h1-hurdle-v0 4 [3, 5] 60 [18, 102] 387 [254, 519] 175 [150, 201] 202 [167, 236]
h1-reach-v0 3203 [2824, 3581] 1409 [998, 1821] 2654 [1951, 3357] 3874 [3220, 4527] 3850 [3272, 4427]
h1-run-v0 4 [2, 6] 91 [54, 128] 778 [763, 793] 232 [185, 279] 415 [307, 524]
h1-stand-v0 15 [7, 22] 433 [138, 727] 798 [779, 817] 772 [701, 843] 814 [770, 857]
h1-walk-v0 8 [1, 16] 33 [22, 45] 814 [813, 815] 550 [391, 709] 845 [840, 850]
IQM 0.007 [0.004, 0.012] 0.134 [0.088, 0.245] 0.734 [0.510, 0.936] 0.521 [0.413, 0.633] 0.799 [0.686, 0.908]
Median 0.021 [0.000, 0.047] 0.284 [0.183, 0.392] 0.696 [0.536, 0.881] 0.598 [0.514, 0.692] 0.781 [0.693, 0.865]
Mean 0.022 [0.000, 0.046] 0.289 [0.207, 0.375] 0.710 [0.562, 0.858] 0.606 [0.536, 0.678] 0.776 [0.705, 0.849]